Jen is 80 years old and her grandson Mark is 5, in how many years will Jen be 4 times Mark's age ?
Answer by macston(5194)

x=years until Jen is 4 times Mark's age
80+x=4(5+x)
80+x=20+4x
60=3x
20=x
ANSWER: Jen will be 4 times Mark's age in 20 years
.
CHECK:
Jen+20=4(Mark+20)
80+20=4(5+20)
100=4(25)
100=100